OpenStudy (anonymous):

A polygon has 11 sides. What is the sum of the measure of the interior angles of the polygon? 1800° 1620° 1440° 1980°

OpenStudy (anonymous):

(n-2) * 180 where n is the number of sides. (11-2)*180 =9*180 = 1620. Good luck

OpenStudy (anonymous):

Convex. (simple) polygon